OpenStudy (anonymous):

A jar has 5 marbles, 1 of each of the colors red, white, blue, green, and yellow. If 4 marbles are removed from the jar, what is the probability that the yellow marble was removed?

OpenStudy (anonymous):

i remember this, yu gotta use factorial

OpenStudy (anonymous):

what is the probability that the marble is NOT removed. If you know that then you can subtract that from 1 to get the answer you seek

OpenStudy (anonymous):

first is not yellow \(\frac{4}{5}\) second not yellow given first isn't \(\frac{3}{4}\) this is not yellow given the first two is \(\frac{2}{3}\) and finally 4th is not yellow is \(\frac{1}{2}\) multiply these together, subtract from 1

OpenStudy (anonymous):

@jon.stromer.galley thx for helping too!

OpenStudy (anonymous):

@satellite73 problem solved thank you!